Fermi National Laboratory
 

Operation Summary for July 6, 2011 - July 8, 2011

 

Wednesday July 6

The day shift began with Operations monitoring the Tevatron (TEV) store 8862 (58.1E30) and the Recycler stash (267.6E10), with the Antiproton Source (Pbar) stacking (9.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Kicker personnel reported at 8:16 AM that they began repairs on the MI kicker, I:KPS5S. Operations prepared Pbar for access and safety system tests at 9:03 AM. Operations established beam to Meson FTBF T-1015 experiment at 9:16 AM. Operations terminated store 8862 at 9:43 AM. Operators prepared D0 and CDF for accesses at 9:58 AM. A Booster RF station (BRF3) tripped off at 10:50 AM for the first of seven times in five hours; it reset. Technicians replaced a faulty relay for the TEV sector D3 LCW pump. Pbar safety system testing ended at 2:55 PM.

During the evening shift, Operators put the TEV through the first of three HEP ramps at 4:41 PM. Operators put the TEV through a dry squeeze at 4:57 PM. Operations disabled beam to T-1015 at 6:19 PM. A Kicker person reported at 7:09 PM that KPS5S was ready. Operations performed several pilot transfers to validate the kicker timing adjustments. TEV proton injection closure looked good. Operators put the TEV through the first of three wet squeezes at 7:55 PM. Operations put the TEV into shot setup at 9:31 PM. At 10:16 PM, KPS5S tripped off; it reset. Operations established store 8866 at 23:02 PM with an average initial luminosity (AIL) of 335.27E30. At 11:22 PM, a TEV kicker (T:E48K) tripped off; it will be reset between stores.

Thursday July 7

The midnight shift began with Operations monitoring store 8866 (235.46E30) and the stash (70.1E10), with Pbar stacking (10.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 1:35 AM, the H- Source high voltage tripped off; it reset. Operations established beam to T-1015 at 4:56 AM.

During the day shift, FESS and Mechanical Support personnel began work on the TEV sector C1 pond water strainer at 8:52 AM; they completed the job in 30 minutes and moved on to D1 and then to D4. Kicker personnel replaced the Glassman High Voltage power supply for the TEV E48 kicker. A TEV RF station (TRF1) tripped off with a downstream heater flow interlock indication; this caused a luminosity loss of 5E30. Operations established beam to MTA at 2:34 PM. Operations terminated store 8866 at 3:10 PM. A TEV RF person repaired TRF1’s flow meter and brought the station online. Operations disabled the MTA beam at 3:38 PM. Operations put the TEV into shot setup at 3:46 PM. FESS personnel reported at 3:51 PM that the TEV C3 pond water pump was not operating properly; this will be investigated further on Friday (7/8/11).

During the evening shift, at 4:03 PM TRF1 tripped off again; RF personnel repaired it and brought it online at 4:42 PM. At 4:58 PM, operators reported that the second injection of antiprotons went into the wrong bucket. Operators investigated, conferred with TEV personnel and the Run Coordinator, and halted shot setup. Operators put the TEV through a dry squeeze. Operations put the TEV into shot setup at 6:09 PM and established store 8868 at 7:12 PM with an AIL of 380.02E30. Operations disabled beam to T-1015 at 8:06 PM. At 10:46 PM, the Recycler stash (112E10) was lost due to the LCW pumps at MI-60 tripping off, which caused Recycler corrector magnet power supplies to trip off. Operators reported short downtimes due to problems with LRF1 & 5, BRF16, MIRF14, and a MIRF South Anode trip.

Friday July 8

The midnight shift began with Operations monitoring store 8868 (154.6E30) and the Pbar stack (13.3E10), and with NuMI and MiniBooNE not taking beam due to the MI-60 LCW pump problems.

During the midnight shift, MI LCW problems continued. LRF5 tripped off at 12:43 AM; it reset. Operators reported at 12:47 AM that the MiniBooNE special LCW water skid had tripped off when the MI-60 LCW tripped off. Operators brought the Pelletron back on line at 12:49 AM. Operators and a MI Specialist reported that the MI-60 LCW pump #3 was the only pump that had a breaker trip. They consulted with the MI Coordinator and then resumed normal operations. Beam to NuMI and Pbar resumed at 1:07 AM. Operators contacted MiniBooNE and Safety personnel, reset the MiniBooNE water skid, and resumed sending beam at 1:29 AM. Operations established beam to T-1015 at 5:03 AM. The MI-10 upper house bypass interlock loop tripped off; it reset.

The Plans for Friday and the Weekend

The plans for today are to terminate store 8868 around noon and have operators prepare the TEV for access. The Run Coordinator will allow Vacuum technicians to work in the TEV, CDF personnel to access their detector, and TEV Kicker personnel to investigate problems with the E48 kicker. He will also allow NuMI personnel to conduct a target scan, Booster personnel to conduct studies, and FESS personnel to conduct thermal scans of the MI LCW breaker panels and pump motors.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/08/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory